Circular No. 6263

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The Supernova Cosmology Project [S. Perlmutter, S. Deustua, S. Gabi,

G. Goldhaber, A. Goobar (now Univ. of Stockholm), D. Groom, B. Grossan, I. Hook, A. Kim, M. Kim, J. Lee, R. Pain (CNRS, Paris), C. Pennypacker, and I. Small (Lawrence Berkeley Lab. and the Center for Particle Astrophysics, Berkeley); R. Ellis and R. McMahon, Inst. of Astronomy, Cambridge; and B. Boyle, P. Bunclark, D. Carter, and M. Irwin, Royal Greenwich Obs.; with K. Glazebrook (Anglo-Australian Obs.) at the W. Herschel Telescope (WHT); A. V. Filippenko and T. Matheson (Univ. of California, Berkeley) at the Keck telescope; W. Couch (Univ. of N.S.W.); M. Dopita and J. Mould (Mt. Stromlo and Siding Spring Obs.) at the Siding Spring 2.3-m telescope; H. Newberg (Fermi National Accelerator Lab.) at the ARC telescope; M. Postman and W. Oegerle (Space Telescope Science Inst.), T. Lauer (National Optical Astronomy Obs.), and J. Hoessel (Univ. of Wisconsin) at the Kitt Peak 4-m telescope] reports the discovery of four additional supernovae in the course of the 1993–94 High Redshift Supernova Search (cf. IAUC 5956):

SN  1993/94  α_{2000}  δ_{2000}  R  Offset
1993al Dec. 22 12^{h}22^{m}12.^{s}62 − 1^{o}08′24′′.5 22.6 2′′.0 W, 0′′.3 S
1994al Jan. 8 3 06 22.45 +17 18 33.7 22.6 0′′.9 E, 0′′.1 N
1994am Feb. 4 2 40 02.06 − 1 37 14.9 21.7 0′′.6 W, 0′′.9 S
1994an Oct. 31 22 44 18.87 + 0 06 48.7 22.3 0′′.2 E, 0′′.2 N

All but SN 1994am (Kitt Peak 4-m reflector) were discovered with the Isaac Newton Telescope. The date and telescope for acquisition of the host-galaxy spectrum, galaxy redshift, approximate time from maximum light at discovery (in days), and date of previous recorded image of galaxy not showing each supernova (to limiting mag ∼ 24) are as follows: SN 1993al, 1994 Jan. 14, Keck (T. Bida, J. Cohen, and J. B. Oke), 0.18 (probable host), —, 1992 Apr. 25; SN 1994al, 1994 Nov. 9, Keck, 0.420, −6, 1993 Dec. 19; SN 1994am, 1994 Nov. 7, WHT, 0.372, 0, 1993 Dec. 19; SN 1994an, 1994 Nov. 9, Keck, 0.378, −1, 1994 Oct. 9. SNe 1993al, 1994al, and 1994am were observed during 1994 Jan. 10–Feb. 16, and SN 1994an during Oct. 31–Dec.

6. The photometry for SNe 1994al, 1994am, and 1994an are all consistent with type-Ia supernovae at the host-galaxy redshift. The spectrum of SN 1994an on 1994 Nov. 9 was identified as a type-Ia SN, ≈ 7 days past maximum light (rest frame) at the host-galaxy redshift.

1995 November 18 Daniel W. E. Green